# A combined Quantum Monte Carlo and DFT study of the strain response and magnetic properties of two-dimensional (2D) 1T-VSe2 with charge density wave

### v1

@@ III Results and Discussion
Fig. 3 illustrates the magnetic phase transitions for both (a) undistorted and (b) charge density wave (CDW) phases. In each case, the exchange interaction is predominant over other interactions. Notably, is over three times higher in the undistorted phase, measuring 71 meV, compared to 21 meV in the CDW phase. This pronounced disparity in interaction strength indicates that spins in the undistorted structure align more strongly with neighbors, leading to an extended correlation length and larger clusters of aligned spins, requiring more thermal energy to disrupt these interactions and disorder the aligned spins. This implies that there exists two potential magnetic phase transitions for freestanding (no substrate, strain or defects) 1T-VSe2, one for the undistorted monolayer below room temperature ( 100 K less than the transition temperature reported in Ref. 9) and one for the CDW-distorted (slightly below liquid N2). These results indicate that the magnetic phase transition occurs below the CDW transition temperature, which has been reported to be above 121 K. This is expected since the CDW phase must exist before the CDW-FM phase can exist by further cooling. The small energy scale between the CDW-distorted phases and the undistorted phases (reported in our current work and previous work[26]) poses a challenge in understanding which phase is energetically more favorable. This small energy scale opens the door to tune the energetic stability of various magnetic states (with and without CDW-distortion) through external mechanisms such as strain, defects, substrate engineering or temperature.
@@
From Fig. 4a), we observe that the CDW-distorted structure in FM and AFM configurations is more sensitive to small amounts of strain in contrast to the undistorted structure. In other words, the energy differences between the strain curves are much smaller for the CDW-distorted structures (compared to undistorted). Notably, we observe that for positive values of strain (lattice expansion), the FM in the undistorted and CDW-distorted structures becomes more stable. For the undistorted structure, we see that the energy difference between the FM state and AFM state drastically increases with the application of positive (tensile) strain. We observe that with small amounts of compressive strain, the undistorted FM structure becomes more favorable than the CDW-distorted FM structure, which can be a promising route to stabalizing the FM in the undistorted structure. For larger amounts of compressive strain ( ), AFM order begins to become more favorable for the undistorted and CDW-distorted (specifically AFMB1) structures.
@@
To further understand this phenomena, we computed under strain for the energy differences depicted in Fig. 4a) for the undistorted and CDW-distorted structures. We then went on to estimate under strain using these values of for additional Monte Carlo simulations (similar to those reported in Fig. 3). It is important to note that for these Monte Carlo simulations under lattice strain, we did not include and in the Hamiltonian (due to increased computational demand of spin-orbit SCAN calculations). Since the driving force behind the is , we do not expect major fluctuations in from and with applied strain. These results are depicted in Fig. 4b). In addition to the strong competition between the undistorted FM and the CDW-FM structures (dotted purple and orange curves in Fig. 4a)), there is potential strain-tunability within each FM ground state (with and without CDW). In Fig. 4b), we see that can be enhanced in the CDW-FM state by small amounts of biaxial expansion or compression, with the approaching values above 120 K. With regards to the undistorted FM structure, we find that tensile strain can enhance the to be above room temperature, but there is strong competition with the FM-CDW state for small values of tensile strain, with the CDW-FM becoming significantly more energetically favorable after strain. In addition to the tunability of magnetic states, it was revealed from anharmonic phonon calculations that the CDW order can be tuned with small amounts of lattice strain (as little as 1.5 ), highlighting the strong substrate dependence [81]. For example, common substrates for VSe2 [9] such as MoS2 (a = b = 3.16 Å) and highly oriented pyrolytic graphite (a = b = 2.46 Å) have significant lattice mismatch, which can induce strain (since the lattice constant of monolayer 1T-VSe2 is estimated to be 3.41 Å[8]). As an alternative to substrate engineering, substitutional doping can be a promising route to inducing tensile strain in 1T-VSe2, which can stabalize and even enhance the magnetic properties in both the undistorted and CDW-distorted phases.
300 K | Bulk FM | ML FM | ML AFM | NM |
A1g | 199 | 197 | 197, 139 | 208 |
Eg | 137 | 135 | 130 | |
Bg | 133 | |||
CDW-ML | NM | FM | ||
@@
To validate our DMC-informed DFT calculations of the Raman modes, we grew a VSe2 crystal flake by Chemical Vapor Transport (CVT) (see SI for details). The 1T phase of VSe2 was confirmed by X-Ray diffraction (XRD) (see Fig. S1). Thin flakes of 1T-VSe2 were exfoliated and encapsulated in h-BN (Fig. S2(a)) and subsequent Raman spectra measurements were performed. The Raman spectra were averaged (30 seconds, 20 accumulations) at four distinct positions on a 1T-VSe2 flake capped with h-BN. Fig. 5(a) shows an optical image with the approximate locations where the spectra were measured and Fig. 5(b) shows the corresponding spectra. The thinnest region measured to be about 1.7 nm by atomic force microscope (AFM) Figure S2, position 3, was used for comparison to DFT predictions. To assign the phonon modes polarized Raman spectra were collected from the 1.7 nm thick region (position 3). Data is shown in both the parallel and cross configuration, i.e., the incoming laser excitation and the Raman scatter were either parallel to one another, or at 90-degrees, confirming the respective E and A symmetry of the modes of the VSe2. In the 1.7 nm thick region we identify two modes; the A1g at 206.7 cm-1 0.4 and the Eg at 137.5 cm-1 0.3 in good agreement with the two modes predicted by DFT. The other features correspond to the Si/SiO2 substrate, as can be seen when compared to measurement of the blank substrate (position 4). However, in the bulk-like region on the h-BN capped flake, position 1, there are additional observable features to the right of the A1g mode which differ from the edge-like feature observed in position 3 coming from the substrate. Previously a broad mode at about 257 cm-1 has been attributed as an Eg mode of the 1T phase in a bulk sample [82], however, we observe in some cases two distinct modes at 235.3 cm-1 and 253.2 cm-1 in the bulk-like region in the h-BN capped flake and at 229.8 cm-1 and 250.3 cm-1 in the bulk crystal, shown in Figure S3. Since these are only observed in the bulk-like capped region/crystal, and are not predicted by the DFT results, these modes are likely due Se [83, 84, 85]. We conclude that the 1T phase of VSe2 only has two phonon modes which we identify as the higher intensity A1g and the lower intensity Eg.
IV Conclusion
@@
We have provided a comprehensive investigation of the magnetic properties and strain response of monolayer 1T-VSe2 using DFT and highly accurate DMC methods. By extensively benchmarking various DFT functionals against our DMC results, we demonstrated the significant impact of the exchange-correlation functional on the accuracy of magnetic properties and energetic stability of various magnetic states of 2D 1T-VSe2, with and without CDW. Our high-fidelity results aim to resolve previous discrepancies that exist in the theoretical and experimental literature for 1T-VSe2. Through classical Monte Carlo simulations informed by DMC and DFT, we estimated the magnetic transition temperatures for the undistorted (228 K) and CDW-distorted (68 K) phases and our strain calculations indicated that small amounts of biaxial strain can enhance the transition temperature, which provides a viable route for engineering the magnetic properties. Furthermore, our Raman spectroscopy experiments on exfoliated flakes of 1T-VSe2 validate our predictions and provide insight on the intrinsic strain mechanisms. Our work underscores the important role of highly accurate many-body methods such as DMC in describing the electronic and magnetic properties of 2D materials. We hope our combined approach can pave the way for future explorations of correlated 2D magnetic materials.
V Notes
@@ VI Competing interests
VII acknowledgments
This work was supported by the National Science Foundation through the Division of Materials Research under NSF Grant No. DMR-2213398. The authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. Work by N.G. was supported by the National Institute of Standards and Technology Summer Undergraduate Research Fellowship program. Work by K.S. and J.T.K. (discussion, analysis of QMC calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Work by T. B. (discussion, analysis of DFT+U calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. F.A. and B.D. would like to acknowledge the support of Mr. Peter Morrison and the Office of Naval Research under Contract No. N00014-21-WX-01248. Research at the United States Naval Academy (M.E.J.) was supported by the Office of Naval Research under Contract No. N0001423WX02132.

### v2

@@ II Results and Discussion
Fig. 2 illustrates the magnetic phase transitions for both (a) undistorted and (b) CDW phases. In each case, the exchange interaction is predominant over other interactions. Notably, is over three times higher in the undistorted phase, measuring 71 meV, compared to 21 meV in the CDW phase. This pronounced disparity in interaction strength indicates that spins in the undistorted structure align more strongly with neighbors, leading to an extended correlation length and larger clusters of aligned spins, requiring more thermal energy to disrupt these interactions and disorder the aligned spins. This implies that there exists two potential magnetic phase transitions for freestanding (no substrate, strain or defects) 1T-VSe2, one for the undistorted monolayer below room temperature ( 100 K less than the transition temperature reported in Ref. 9) and one for the CDW-distorted (slightly below liquid N2). These results indicate that the magnetic phase transition occurs below the CDW transition temperature, which has been reported to be above 121 K. This is expected since the CDW phase must exist before the CDW-FM phase can exist by further cooling. The small energy scale between the CDW-distorted phases and the undistorted phases (reported in our current work and previous work[26]) poses a challenge in understanding which phase is energetically more favorable. This small energy scale opens the door to tune the energetic stability of various magnetic states (with and without CDW-distortion) through external mechanisms such as strain, defects, substrate engineering or temperature.
@@
From Fig. 3a), we observe that the CDW-distorted structure in FM and AFM configurations is more sensitive to small amounts of strain in contrast to the undistorted structure. In other words, the energy differences between the strain curves are much smaller for the CDW-distorted structures (compared to undistorted). Notably, we observe that for positive values of strain (lattice expansion), the FM in the undistorted and CDW-distorted structures becomes more stable. For the undistorted structure, we see that the energy difference between the FM state and AFM state drastically increases with the application of positive (tensile) strain. We observe that with small amounts of compressive strain, the undistorted FM structure becomes more favorable than the CDW-distorted FM structure, which can be a promising route to stabilizing the FM in the undistorted structure. For larger amounts of compressive strain ( ), AFM order begins to become more favorable for the undistorted and CDW-distorted (specifically AFM-B1) structures.
@@
To further understand this phenomena, we computed under strain for the energy differences depicted in Fig. 3a) for the undistorted and CDW-distorted structures. We then went on to estimate under strain using these values of for additional Monte Carlo simulations (similar to those reported in Fig. 2). It is important to note that for these Monte Carlo simulations under lattice strain, we did not include and in the Hamiltonian (due to increased computational demand of spin-orbit SCAN calculations). Since the driving force behind the is , we do not expect major fluctuations in from and with applied strain. These results are depicted in Fig. 3b). In addition to the strong competition between the undistorted FM and the CDW-FM structures (dotted purple and orange curves in Fig. 3a)), there is potential strain-tunability within each FM ground state (with and without CDW). In Fig. 3b), we see that can be enhanced in the CDW-FM state by small amounts of biaxial expansion or compression, with the approaching values above 120 K. With regards to the undistorted FM structure, we find that tensile strain can enhance the to be above room temperature, but there is strong competition with the FM-CDW state for small values of tensile strain, with the CDW-FM becoming significantly more energetically favorable after strain. In addition to the tunability of magnetic states, it was revealed from anharmonic phonon calculations that the CDW order can be tuned with small amounts of lattice strain (as little as 1.5 ), highlighting the strong substrate dependence [56]. For example, common substrates for VSe2 [9] such as MoS2 (a = b = 3.16 Å) and highly oriented pyrolytic graphite (a = b = 2.46 Å) have significant lattice mismatch, which can induce strain (since the lattice constant of monolayer 1T-VSe2 is estimated to be 3.41 Å[8]). As an alternative to substrate engineering, substitutional doping can be a promising route to inducing tensile strain in 1T-VSe2, which can stabilize and even enhance the magnetic properties in both the undistorted and CDW-distorted phases.
| Bulk FM | ML-FM | ML-AFM | ML-NM | Exp | |
| A1g | 199 | 197 | 197, 139 | 208 | 206.7(4) |
| Eg | 137 | 135 | 130 | 137.5(3) | |
| Bg | 133 | ||||
| CDW-ML | NM | FM | |||
@@
To validate our DMC-informed DFT calculations of the Raman modes, we grew a VSe2 crystal by Chemical Vapor Transport (CVT) (see SI for details). The 1T phase of VSe2 was confirmed by X-Ray diffraction (XRD) (see Fig. S2). Thin flakes of 1T-VSe2 were exfoliated and encapsulated in h-BN (Fig. S3(a)) and subsequent Raman spectra measurements were performed. The Raman spectra were averaged (30 seconds, 20 accumulations) at four distinct positions on a 1T-VSe2 flake capped with h-BN. Figure 4(a) shows an optical image with the approximate locations where the spectra were measured and Fig. 4(b) shows the corresponding spectra. The thinnest region measured to be about 1.7 nm by atomic force microscope Figure S3, position 3, was used for comparison to DFT predictions. To assign the phonon modes polarized Raman spectra were collected from the 1.7 nm thick region (position 3). Data is shown in both the parallel and cross configuration, i.e., the incoming laser excitation and the Raman scatter were either parallel to one another, or at 90-degrees (cross), confirming the respective A and E symmetry of the modes of the VSe2. In the 1.7 nm thick region we identify two modes; the A1g at 206.7 cm-1 0.4 and the Eg at 137.5 cm-1 0.3 in good agreement with the two modes predicted by DFT. The other features correspond to the Si/SiO2 substrate, as can be seen when compared to measurement of the bare substrate (position 4). However, in the bulk-like region on the h-BN capped flake, position 1, there are additional observable features to the right of the A1g mode which differ from the edge-like feature observed in position 3 coming from the substrate. Previously a broad mode at about 257 cm-1 has been attributed as an Eg mode of the 1T phase in a bulk sample [57], however, we observe in some cases two distinct modes at 235.3 cm-1 and 253.2 cm-1 in the bulk-like region in the h-BN capped flake and at 229.8 cm-1 and 250.3 cm-1 in the bulk crystal, shown in Figure S4. Since these are only observed in the bulk-like capped region/crystal, and are not predicted by the DFT results, these modes are likely due Se [58, 59, 60]. We conclude that the 1T phase of VSe2 only has two phonon modes which we identify as the higher intensity A1g and the lower intensity Eg.
III Conclusion
@@
We have provided a comprehensive investigation of the magnetic properties and strain response of monolayer 1T-VSe2 using DFT and highly accurate DMC methods. By extensively benchmarking various DFT functionals against our DMC results, we demonstrated the significant impact of the exchange-correlation functional on the accuracy of magnetic properties and energetic stability of various magnetic states of 2D 1T-VSe2, with and without CDW. Our high-fidelity results aim to resolve previous discrepancies that exist in the theoretical and experimental literature for 1T-VSe2. Through classical Monte Carlo simulations informed by DMC and DFT, we estimated the magnetic transition temperatures for the undistorted (228 K) and CDW-distorted (68 K) phases and our strain calculations indicated that small amounts of biaxial strain can enhance the transition temperature, which provides a viable route for engineering the magnetic properties. Furthermore, our Raman spectroscopy experiments on exfoliated flakes of 1T-VSe2 validate our predictions. Our work underscores the important role of highly accurate many-body methods such as DMC in describing the electronic and magnetic properties of 2D materials. We hope our combined approach can pave the way for future explorations of correlated 2D magnetic materials.
IV Methods
Benchmarking DFT calculations for monolayer 1T-VSe2 were performed with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials [61, 62]. For testing purposes, we employed a variety of exchange-correlation functionals including the local density approximation (LDA)[63, 64], Perdew-Burke-Ernzerhof (PBE)[65], the strongly constrained and appropriately normed (SCAN)[66] meta-GGA functional and the r2SCAN [50] meta-GGA functional with, and without the Hubbard correction [33], which was used to treat the on-site Coulomb interaction of orbitals of V. A k-grid of was used and scaled proportionally to the supercell size, a plane-wave kinetic energy of 400 eV (increased to 800 eV for SCAN and r2SCAN) was used and at least 20 Å of vacuum was added between periodic layers. Our density functional perturbation theory (DFPT) [67] phonon calculations to obtain the peak positions of Raman modes were also performed with VASP along with the phonopy [68, 69] package.
Our Quantum Monte Carlo (QMC) calculations for 1T-VSe2 used the exact same settings as our previous work on 1T- and 2H-VSe2 (see Ref. 8 for more specific details regarding kinetic energy cutoff, k-point grid, finite-size, and timestep convergence). DFT-PBE was used to generate the trial wavefunction (nodal surface) for DMC using the Quantum Espresso (QE) [70] code with a correction of 2 eV. This was due to the fact that PBE+ = 2 eV resulted in a trial wavefunction with the lowest total energy calculated with DMC [8]. We used a k-point grid of and a kinetic energy cutoff of 4,080 eV (300 Ry) to generate the trial wavefunction. In terms of pseudopotentials, RRKJ (OPT) potentials were used for V [71] and Burkatzki-Fillipi-Dolg (BFD) were used for Se [72]. Variational Monte Carlo (VMC) and DMC [32] simulations were conducted using the QMCPACK [73, 74] code and the Nexus [75] workflow manager was used to automate the DFT-VMC-DMC calculations. Up to two-body Jastrow parameters [76, 77, 78] were optimized using the linear method [79, 80] to minimize the variance and energy in VMC, with the goal of modeling electron correlation and reducing uncertainty in the DMC results [80, 81]. In order to compute the nonlocal part of the pseudopotentials, the locality approximation [81] was used in DMC. An optimal timestep of 0.01 Ha-1 (0.27 eV-1) [8] was used in DMC and to reduce finite-size errors, canonical twist averaging (with uniform weighting) [82] and extrapolating to larger supercells (up to 72 atoms) was performed. A more detailed explanation of the theory behind VMC and DMC can be found in Ref. 32.
Energies of different spin configurations derived from first-principles calculations were mapped onto a classical spin Hamiltonian, characterizing three-dimensional spins arranged on a two-dimensional lattice at sites of V atom. Subsequently, this Hamiltonian served as the basis for executing classical Monte Carlo simulations aimed at estimating the transition temperature () for both the undistorted and CDW phases. The Hamiltonian, which has previously been used for other 2D magnetic systems [83, 84, 37], includes the bilinear isotropic exchange (), anisotropic exchange (), and easy axis single ion anisotropy (),
The indices and iterate over all magnetic sites and their corresponding first nearest neighbor (NN) magnetic sites, respectively. Only the first NN interactions are considered due to the strong localized magnetic moment on V. We have calculated for the CDW phase to be , supporting these assumptions (using SCAN). According to the adopted sign convention, favors isotropic spin alignment, leading to a preferred ferromagnetic (FM) over an antiferromagnetic (AFM) phase. Similarly, favors alignment of the spin -components, and favors the out-of-plane direction as the easy axis. We employ a supercell, equivalent to a nearly nm2 square cell, for the Monte Carlo (MC) simulations. The spin configuration space is discretized into uniformly distributed points on the surface of a unit sphere with a fine resolution of degrees, facilitating detailed exploration of spin orientations. The MC simulation consists of numerous sweeps across all magnetic sites. During each sweep, the spin orientation at each site is updated randomly, one site at a time, and the new configuration is then either accepted or rejected based on the Metropolis algorithm. The convergence of both energy and magnetization is meticulously monitored across varying temperatures. Upon achieving equilibration, an ensemble of configurations is collected to calculate the magnetization and specific heat. We use an ensemble size of thousand for the CDW simulation and thousand for the undistorted simulation. Additional methodology details can be found in the SI.
V Supporting Information
Additional discussion of DFT benchmarking for undistorted and CDW structures, theoretical details of Heisenberg spin Hamiltonian and extraction of magnetic parameters, additional details of experimental synthesis and characterization
VI Data Availability Statement
The data from the present work is available on Figshare via https://doi.org/10.6084/m9.figshare.28127477.
VII Code Availability Statement
The code used to perform classical Monte Carlo simulations using the spin Hamiltonian is available at https://github.com/UMBC-STEAM-LAB/SpinMCPack.
VIII Notes
@@ IX Competing interests
X acknowledgments
This work was supported by the National Science Foundation through the Division of Materials Research under NSF Grant No. DMR-2213398 and the Department of Energy (DOE) under Grant DE-SC0024236. The authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. Work by N.G. was supported by the National Institute of Standards and Technology Summer Undergraduate Research Fellowship program. Work by K.S. and J.T.K. (discussion, analysis of QMC calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Work by T. B. (discussion, analysis of DFT+U calculations) was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. F.A. and B.D. would like to acknowledge the support of Mr. Peter Morrison and the Office of Naval Research under Contract No. N00014-21-WX-01248. Research at the United States Naval Academy (M.E.J.) was supported by the Office of Naval Research under Contract No. N0001423WX02132.
Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).